1.    The volume sample of gas is 750 mL at 60°C and 0.7 atm. At what temperature will the sample 1.75 L occupy under a pressure of 1.80 atm?


1
Expert's answer
2022-04-17T11:04:24-0400

Ideal gas law: "p*V=n*R*T"

Where p is pressure, V is volume, n is the amount of gas, R is ideal gas constant and T is temperature


R and n remain constant, which means "\dfrac{p*V}{T}=constant"

Hence, "\dfrac{p_1*V_1}{T_1}=\dfrac{p_2*V_2}{T_2}"


"p_1 =0.7\ atm,\ p_2 =\ 1.80\ atm"

"V_1=750\ ml,\ V_2=1750\ ml"

"T_1 =60^oC=60+273.15=333.15K"


"\dfrac{0.7*750}{333.15}=\dfrac{1.80*1750}{T_2}" ,


"T_2 = 1998.9K=1998.9-273.15=1725.75^oC"